How do you solve
5/6 / a/b = 1/4
trying to find the value of a/b

Answer by rfer(16322)

5/6=1/4*a/b
(5/6)/(1/4)=a/b
5/6*4/1=a/b
20/6=a/b
3 1/3=a/b